The price elasticity of gasoline supply in the U.S is 0.4. If the price of gasoline rises by 8% , what is the expected change in the quantity of gasoline supplied in the U.S?
1
Expert's answer
2020-05-08T13:16:19-0400

1.By definition price elasticity of supply is % change in quantity supplied divided by % change in  price:

"E=\frac{\Delta Q}{\Delta P}"

where "\Delta Q" - change in quantity supplied measured in percents,

"\Delta P" - change in price measured in percents.


2.This means that expected change in quantity of gasoline supplied in U.S.can be calculated as:

"\Delta Q=E*{\Delta P} =0.4*8=3.2" %


Quantity of gasoline supplied in U.S.is expected to increse by 3.2%.